OpenStudy (anonymous):

Find the rational roots of x4 +8x3 +7x2 −40x−60 = 0

OpenStudy (anonymous):

Try to factor x4 +8x3 +7x2 −40x−60

OpenStudy (anonymous):

Factor theorem may help.

OpenStudy (shubhamsrg):

well,,you need to find only rational roots..rational root theorem is what you are looking for..

OpenStudy (anonymous):

x^4 +8*x^3 +7*x^2 -40*x -60 = 0 => => x^4 +8*x^3 +7*x^2 -40*x -60 = 0 => (x +6)*(x +2.236)*(x +2)*(x -2.236) = 0 Real solutions: Root 1: -6 Root 2: -2.236 Root 3: -2 Root 4: 2.236